OpenStudy (anonymous):

Can someone help me with geometry! i am stressed out reaching out to higher educated people!

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

what are the center and radius of a circle with equation (x+2)^2 +(y-9)^2=16? center :(-9,2) r=4 center:(-2,9)r=16 Center :(2,-9) r=16 center (-2,9) r=4

OpenStudy (anonymous):

The equation of a circle is x^2 + y^2 = r^2. That circle is centered on the origin. To center it on a point (j, k) you subtract the coordinates. (x - j)^2 + (y - k)^2 = r^2 That is called translating. You have it easy: you are given the equation in that form, so you don't even have to find factors. Just read the answers out of the equation. The radius is 4, the center is (-2, 9).

OpenStudy (anonymous):

so D?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sweet thanks

OpenStudy (anonymous):

can you help me with 3 more?

OpenStudy (anonymous):

um....sure...ill try

OpenStudy (anonymous):

ok :) sorry i SUCK at math

OpenStudy (anonymous):

its ok...

OpenStudy (anonymous):

what is the equation of the circle with center (0,0) that passes through the point (-6,-6)? (x-(-6))^2+(y-(-6))^2=72 x^2+y^2=0 x^2+y^2+72 (x-(-6))^2 +(y-(-6))^2=0

OpenStudy (anonymous):

@MeganChase

OpenStudy (anonymous):

um...im working it...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the last one maybe

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

WZ and XR are diameters. What is the measure of ZWX? 214 266 60 300 I had the most difficulty with this one. lol

OpenStudy (anonymous):

you would do 180-86-34=60

OpenStudy (anonymous):

for these is it always 180?

OpenStudy (anonymous):

last question

OpenStudy (anonymous):



OpenStudy (anonymous):

no the diameters cuts the circle in half which it equals 180

OpenStudy (anonymous):

If m AB=50 degrees and M CD=24 degrees, what is the value of x? x=26 x=62 x=37 x=74

OpenStudy (anonymous):

OHHH that makes since. the full circle is 360 and the diameters would automaticly make it 180.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i got 26 for the last question

OpenStudy (anonymous):

umm...hold on the last question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh wait you got 26 for the question you just put up?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh...then yes...thats what i got...

OpenStudy (anonymous):

ok SWEET THANKS SO MUCH! :D :D

OpenStudy (anonymous):

yep!